A 280-kg object and a 580-kg object are separated by 3.50 m.

(a) Find the magnitude of the net gravitational force exerted by these objects on a 33.0-kg object placed midway between them.
_________N

(b) At what position (other than an infinitely remote one) can the 33.0-kg object be placed so as to experience a net force of zero from the other two objects?
_____________m from the 580 kg mass toward the 280 kg mass
Answer #1

m = 2804 2 m2 = 58kg 7075m Fnet = F2-F F iumem cum a umam amim S dd = 6.67X10 x 33 580-280 10752 Fnet = 2.15 xlotu d=3-som m